OpenStudy (anonymous):

Find the equations of two different circles so that each circle satisfies the conditions if the center is on the x-axis and graph is tangent to the horizontal line through the point (0,3)

OpenStudy (anonymous):



OpenStudy (anonymous):

equation of a circle with radius\(r\) and center \((h,k)\) is \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

radius is evidently 3 so you have \[(x-h)^2+(y-k)^2=9\] since the center is on the \(x\) axis, the \(y\) coordinate of the center is \(0\) so it is going to look like \[(x-h)^2+y^2=9\] to make up two examples, pick a couple different values for \(h\)